A nut store sells cashews for $9 per pound and almonds for $3.50 per pound. At the end of the month, the almonds hadn't sold so well. So the owner decided to mix the almonds with 60 pounds of cashews. The owner wants to sell the mixture for $7.50 per pound. How many pounds of almonds are necessary to obtain the desired mixture that will sell for $7.50 per pound? Write an equation and solve.
Answer by stanbon(75887)

A nut store sells cashews for $9 per pound and almonds for $3.50 per pound. At the end of the month, the almonds hadn't sold so well. So the owner decided to mix the almonds with 60 pounds of cashews. The owner wants to sell the mixture for $7.50 per pound. How many pounds of almonds are necessary to obtain the desired mixture that will sell for $7.50 per pound? Write an equation and solve.
-----
Value Equation:
Value of cahews + Value of Almonds = Value of mixture
9*60 + 3.5a = 7.5*(a + 60)
90*60 + 35a = 75a + 75*60
40a = 15*60
a = (3/2)15
a = 9 (# of lbs. of almonds needed)
==============================
Cheers,
Stan H.
=================